# WISE: Rethinking the Knowledge Memory for Lifelong Model Editing of Large Language Models

### v1

@@ 4 Related Works
Memory and Knowledge Injection of LLMs. LLMs have long-term (episodic) and working memory [22, 23, 25]. Long-term memory is stored in model parameters, updatable via (re)pretraining [51], finetuning [57], and model editing [14]. Working memory resides in neuron activations, utilized during inference [22]. In-context learning and retrieval-based editing methods like GRACE contribute to working memory [58, 10]. However, whether finetuning or retrieval is debated [59, 60]. Also, current knowledge injection methods often suffer from computational overhead [13, 10], catastrophic forgetting [61], and overfitting [62]. Methods like MemorryLLM [26], SPALM [25], NKB [63], and Memoria [23] are proposed to improve the memories from the architecture design perspective.
Model Editing of LLMs. Model editing encompasses constrained finetuning, locating-and-editing, meta-learning, and retrieval-based methods. ROME identifies factual associations and edits efficiently using MLP-based memories [16], extended by MEMIT for mass-editing [17]. T-Patcher adds neurons for edits in LLMs’ feed-forward layers [11]. Meta-learning methods like MEND decouple finetuning gradients to generalize edits [29], complemented by MALMEN addressing cancellation effects [15]. Retrieval-based methods like SERAC and GRACE improve working memory for editing [30, 10]. From single to mass editing and static to lifelong editing, model editing evolves to meet realistic demands. The latest efforts in lifelong editing such as LTE [64], MALMEN [15], and RECIPE [65] require extensive training with domain-specific edits before specific editing, yet we cannot predict the domain of upcoming edits in the editing flow and accessing these data is often impractical or unrealistic. It potentially increases the risks associated with retraining.
@@ A.1 Description of Datasets
The ZsRE question-answering task [44] is extensively studied within the model editing literature [16, 17, 29, 15, 11], where each record contains an editing statement $\mathbf{x}_{e}$, a paraphrase prompt $\mathbf{x}^{\prime}_{e}$, and a locality prompt $\mathbf{x}_{\text{loc}}$. We use the same train/test split as [29] (163196/19086). Notably, only MEND requires fitting a hypernetwork on the training set; other methods discard the training set and perform edits and evaluations on the test set. In practice, we randomly sample 1K and 3K records from the test set to form the edit sets in Section 3.2 and 6.
@@ B.5 Importance of Knowledge Anchor When Merging Models
B.6 Parameter Efficiency
@@ Appendix D Detailed Related Works
Memory and Knowledge Injection of LLMs The memories of LLMs can be divided into long-term (episodic) memory and working memory (short-term) [22, 23, 25]. Long-term memory refers to the knowledge stored in the model’s parameters, which can be updated by (re)pretraining [51], finetuning [57], and model editing [14]. Working memory is stored in sustained activations/representations of neurons, which will be awakened during inference time [22]. In-context learning (ICL) is a kind of working memory [58], also along with retrieval-based editing methods like GRACE [10]. How to reinforce memory and inject/update knowledge for LLMs is a fundamental question [26, 59, 60]. ICL or finetuning? Different works show different conclusions. In [60], the authors find that few-shot finetuning is more generalizable than ICL, especially for out-of-distribution data. In [59], the authors contrast finetuning with retrieval-augmented generation (RAG) in terms of knowledge injection and find that RAG is better in most cases, and combining both will produce the best results. However, finetuning and pretraining are computation-expensive [13, 10] and usually suffer from catastrophic forgetting [61] and overfitting [62]. For ICL and RAG, the working memory is sometimes not controllable, the model may not follow the information of the contexts [22], and the context window is limited [78, 79], and there are works addressing these issues by training controllable ICL [22], long-context techniques [78, 79], and recurrent memory architecture design [26]. SPALM is proposed to add language models with storage modules that resemble both working and long-term memories [25]. Also, model editing is an emerging technology that aims to enable data-efficient, fast, and non-destructive knowledge updates to LLMs [14, 53].
@@
Model Editing of LLMs Model editing can be summarized as the following lines of research. Constrained finetuning: Preliminary model editing uses constrained finetuning to update parameters based on new examples [80, 81]. Locate-and-edit: ROME [16] locates the factual associations in autoregressive LLMs and conducts accurate and efficient edits by taking MLPs as key-value memories. Then, MEMIT [17] extends ROME from single-editing to mass-editing. COMEBA-HK [82] identifies the Local Editing Scope and extends MEMIT for sequential editing. In addition, T-Patcher [11] targets the last feed-forward layer of LLMs, adding an additional neuron for each edit. Meta learning: Recent meta-learning methods use hypernetworks for aiding editing. MEND [29] learns a hypernetwork that can decouple the finetuning gradients into the gradient updates that generalize the edits and won’t damage the performances on unrelated inputs. To remedy the cancellation effect of MEND, MALMEN [15] uses hypernetwork to produce the weight shifts of editing and formulates the weight shift aggregation as the least square problem. Retrieval-based methods: Instead of directly editing the model parameters, retrieval-based methods aim to improve the working memory of LLMs to enable model editing. IKE [83] uses context-edit facts to guide the model when generating edited facts. DeCK [84] employs contrasting knowledge decoding, which enhances the confidence of in-context-based editors in the edited facts. SERAC [30] (a modified version dubbed as DEFER [10]) records edit items in a file and trains additional scope classifier and counterfactual model to detect, retrieve, and generate the edit-related results. Though the editing retriever and generator are neural networks, they are too small to have the power of LLMs. GRACE [10] adopts a discrete codebook of edits for retrieving and replacing the edits’ layer representations during inference. From single editing [16] to mass editing [15, 17], and from static editing to sequential [11] (continual) or lifelong editing [10], model editing is developing to meet more realistic demands.
Model Merging

### v2

@@ 4 Related Works
Memory and Knowledge Injection of LLMs.
LLMs have long-term (episodic) and working memory [24, 25, 27]. Long-term memory is stored in model parameters, updatable via (re)pretraining [53], finetuning [59], and model editing [14]. Working memory resides in neuron activations, utilized during inference [24]. In-context learning and retrieval-based editing methods like GRACE contribute to working memory [60, 10]. However, whether finetuning or retrieval is debated [61, 62]. Also, current knowledge injection methods often suffer from computational overhead [13, 10], catastrophic forgetting [63], and overfitting [64]. Methods like MemorryLLM [28], SPALM [27], NKB [65], and Memoria [25] are proposed to improve the memories from the architecture design perspective.
Model Editing of LLMs.
Model editing encompasses constrained finetuning, locating-and-editing, meta-learning, and retrieval-based methods. ROME identifies factual associations and edits efficiently using MLP-based memories [18], extended by MEMIT for mass-editing [19]. T-Patcher adds neurons for edits in LLMs’ feed-forward layers [11]. Meta-learning methods like MEND decouple finetuning gradients to generalize edits [31], complemented by MALMEN addressing cancellation effects [15]. Retrieval-based methods like SERAC and GRACE improve working memory for editing [32, 10]. From single to mass editing and static to lifelong editing, model editing evolves to meet realistic demands. The latest efforts in lifelong editing such as LTE [66], MALMEN [15], and RECIPE [67] require extensive training with domain-specific edits before specific editing, yet we cannot predict the domain of upcoming edits in the editing flow and accessing these data is often impractical or unrealistic. It potentially increases the risks associated with retraining.
Model Merging
@@ 6 Conclusion
Acknowledgements
We would like to express gratitude to the anonymous reviewers for their kind comments. This work was supported by the National Natural Science Foundation of China (No. 62206246, No. NSFCU23B2055, No. NSFCU19B2027), the Fundamental Research Funds for the Central Universities (226-2023-00138), Zhejiang Provincial Natural Science Foundation of China (No. LGG22F030011), Yongjiang Talent Introduction Programme (2021A-156-G), Information Technology Center and State Key Lab of CAD&CG, Zhejiang University.
@@ A.1 Description of Datasets
The ZsRE question-answering task [46] is extensively studied within the model editing literature [18, 19, 31, 15, 11], where each record contains an editing statement $\mathbf{x}_{e}$, a paraphrase prompt $\mathbf{x}^{\prime}_{e}$, and a locality prompt $\mathbf{x}_{\text{loc}}$. We use the same train/test split as [31] (163196/19086). Notably, only MEND requires fitting a hypernetwork on the training set; other methods discard the training set and perform edits and evaluations on the test set. In practice, we randomly sample 1K and 3K records from the test set to form the edit sets in Section 3.2 and 3.3.
@@ B.5 Importance of Knowledge Anchor When Merging Models
B.6 Ablation Study of Random Prefix Token
Figure 12: Ablation studies on Random Prefix Token (PT) of WISE. Light/Dark colors indicate the Editing Sucess w.o./w. PT addition. ZsRE. LlaMA-2-7B
As described in Section 3.1, we employ random prefix token augmentation to enable the editing knowledge to cope with various contexts. That is, for a single $\mathbf{x}_{e}$, it expands into $(\text{prefix}_{i},\mathbf{x}_{e})$. The prefix is derived from tokens that are randomly generated by the original LM $f_{\Theta}$, serving as an economical data augmentation method. We observe that the editing success rate is compromised (Figure 12). Specifically, for instance, at T=1000, Rel. and Gen. decreased by 0.15 and 0.17, respectively. By utilizing randomly generated prefix tokens, the model is able to learn a broader range of linguistic features, thereby exhibiting greater robustness in practical applications. We believe that access to the "data generator" can deepen the model’s memory of editing samples.
B.7 Parameter Efficiency
@@ Appendix D Detailed Related Works
The memories of LLMs can be divided into long-term (episodic) memory and working memory (short-term) [24, 25, 27]. Long-term memory refers to the knowledge stored in the model’s parameters, which can be updated by (re)pretraining [53], finetuning [59], and model editing [14]. Working memory is stored in sustained activations/representations of neurons, which will be awakened during inference time [24]. In-context learning (ICL) is a kind of working memory [60], also along with retrieval-based editing methods like GRACE [10]. How to reinforce memory and inject/update knowledge for LLMs is a fundamental question [28, 61, 62]. ICL or finetuning? Different works show different conclusions. In [62], the authors find that few-shot finetuning is more generalizable than ICL, especially for out-of-distribution data. In [61], the authors contrast finetuning with retrieval-augmented generation (RAG) in terms of knowledge injection and find that RAG is better in most cases, and combining both will produce the best results. However, finetuning and pretraining are computation-expensive [13, 10] and usually suffer from catastrophic forgetting [63] and overfitting [64]. For ICL and RAG, the working memory is sometimes not controllable, the model may not follow the information of the contexts [24], and the context window is limited [91, 92], and there are works addressing these issues by training controllable ICL [24], long-context  [91, 92], and recurrent memory architecture design [28]. SPALM is proposed to add language models with storage modules that resemble both working and long-term memories [27].
Model Editing of LLMs
@@
Model editing can be summarized as the following lines of research. Constrained finetuning: Preliminary model editing uses constrained finetuning to update parameters based on new examples [93, 94]. Locate-and-edit: ROME [18] locates the factual associations in autoregressive LLMs and conducts accurate and efficient edits by taking MLPs as key-value memories. Then, MEMIT [19] extends ROME from single-editing to mass-editing. COMEBA-HK [95] identifies the Local Editing Scope and extends MEMIT for sequential editing. In addition, T-Patcher [11] targets the last feed-forward layer of LLMs, adding an additional neuron for each edit. Meta learning: Recent meta-learning methods use hypernetworks for aiding editing. MEND [31] learns a hypernetwork that can decouple the finetuning gradients into the gradient updates that generalize the edits and won’t damage the performances on unrelated inputs. To remedy the cancellation effect of MEND, MALMEN [15] uses hypernetwork to produce the weight shifts of editing and formulates the weight shift aggregation as the least square problem. Retrieval-based methods: Instead of directly editing the model parameters, retrieval-based methods aim to improve the working memory of LLMs to enable model editing. IKE [96] uses context-edit facts to guide the model when generating edited facts. DeCK [97] employs contrasting knowledge decoding, which enhances the confidence of in-context-based editors in the edited facts. SERAC [32] (a modified version dubbed as DEFER [10]) records edit items in a file and trains additional scope classifier and counterfactual model to detect, retrieve, and generate the edit-related results. Though the editing retriever and generator are neural networks, they are too small to have the power of LLMs. GRACE [10] adopts a discrete codebook of edits for retrieving and replacing the edits’ layer representations during inference. From single editing [18] to mass editing [15, 19], and from static editing to sequential [11] (continual) or lifelong editing [10], model editing is developing to meet more realistic demands.